OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5. Numerical Answers Expected!

OpenStudy (texaschic101):

3x^2 - y^3 - y^3 - z when x = 3, y = -2 and z = -5 now we sub 3(3^2) - (-2^3) - (-2^3) - (-5) = 3(9) - (-8) - (-8) + 5 = 18 + 8 + 8 + 5 = 39

OpenStudy (texaschic101):

with problems such as these, you have to make sure and pay attention to your signs because one little mistake can mess the whole problem up

OpenStudy (anonymous):

thanks i have to do this big test tomorrow and i didnt understand it tel now